# Dynamic read & write optimizaton with TurtleKV

### v1

@@ 1. Introduction
Key contributions:
  1. 1)
    An exploration of the benefit of trading off memory dynamically for increased write or read performance. The design of an architecture that enables high dynamic tuning range by starting with a unbiased on-disk data structure and dynamically trading memory for performance.
  2. 2)
    TurtleTrees, a novel on-disk data structure that merges the benefits of read-optimized B+-trees with write-optimized Level-tiered LSM-trees. The identification and formal definition of an important class of data structures to which TurtleTrees belong, Bϵ+-trees
  3. 3)
    The concept of checkpoint distance ($\chi$), which allows more batched updates to reside in memory before being written to disk. This limits write amplification as previous batches can expire in memory before being written.
  4. 4)
    TurtleKV, a key/value storage engine that uses TurtleTrees and checkpoint distance to achieve high levels of performance on modern SSDs.
2. Overview
Brodal and Fagerberg’s Lower Bounds (Brodal and Fagerberg, 2003) and the RUM Conjecture (Athanassoulis et al., 2016) show that key/value stores are constrained by a three-way trade-off between main memory usage, read operations, and writes. This means that storage engines have at most two degrees of freedom when selecting a tuning point for a given application: for example, committing to a cost function for reads and memory forces a minimum cost for writes.
A key insight of our work is that, while this trade-off is fundamental, the mechanisms in common use by today’s storage engines to navigate the RUM space are not. This means it is theoretically possible to devise a dictionary type affording mechanisms which trade, for example, writes against memory independent of reads, and reads against memory independent of writes. The goal of our work is to demonstrate mechanisms which do this in a way that avoids coding a given tuning point into the structure of stored data. This allows tuning points to be chosen dynamically, by re-allocating memory in different ways to optimize for the workload at hand, expanding the range of applications which can be supported by a single general-purpose key/value store.
In this section we give a brief review of current practice with respect to trade-off tuning in general purpose key/value storage engines, grouped by tuning goal: read-optimized, write-optimized, and hybrid.. We then present the key ideas of this work: Bϵ+-trees, TurtleTrees, Write Optimization via Checkpoint Distance, and On-Demand Optimization.
2.1. Current Approaches to Trade-off Tuning
Here we describe and critique the major strategies presently in use to navigate the RUM trade-off space based on data structure type.
Figure 2. Bϵ-tree $\rightarrow$ Bϵ+-tree structural transformation
Read-Optimized (B-trees). B-trees (Comer, 1979) provide optimal (Yi, 2012) read performance and can utilize memory to reduce the number of reads required to process queries by caching pages containing tree-nodes. B+-trees, the most commonly implemented B-tree variant, exploit this mechanism by placing key/value data in the leaves, allowing all but the bottom layer of the tree to be cached in main memory so that queries can be answered with a single read.
State-of-the-art approaches to reducing the write cost of updates in B-trees employ the use of a write-ahead log (WAL) to ensure durability of updates, plus a scheme for applying updates to the tree in memory. We refer to these as Log Structured B-trees. Their goal is to defer rewriting pages containing updated keys, in hopes of collecting enough updates per page to amortize I/O. Bw-trees (Levandoski et al., 2013a) do this through the use of delta records applied in-memory to the main tree using lock-/latch-free techniques. The closely related Bf-tree (Hao and Chandramouli, 2024) use an alternate strategy of caching and updating sub-page slices (mini-pages) instead full pages, which has additional benefits in terms of lowering the memory cost of caching ”hot” records.
Bw-trees and Bf-trees achieve good write efficiency either when memory is plentiful or when updates are clustered in a few sub-ranges of the key space. In the general case, however, Log Structured B-trees force a sub-optimal trade-off between memory and write amplification, especially as the total data size $N$ grows with respect to available memory $M$ (Idreos et al., 2019).
Write-Optimized (LSM-trees). Log-Structured Merge (LSM)-trees (O’Neil et al., 1996) are a popular write-optimized alternative to B-trees. By deferring the merge of new updates with existing data, they maximize the amount of effective work achieved by each write operation. Unfortunately, this technique forces read costs (and space requirements) to increase, since it introduces the possibility that updates for a given key may be stored in multiple locations. Page caching can be used to trade memory against reads, but this mechanism is less effective for LSM-trees than B-trees because there is potentially much more data to cache.
In addition to caching, two primary tuning mechanisms are employed by LSM-Tree systems: compaction policy and approximate memory query (AMQ) filtering. Compaction policies divide into greedy (level-tiering) and lazy (size-tiering). In a level-tiered LSM-Tree, write cost is increased by repeatedly merging data at the same level (up to $F$ times) to reduce future read cost. In a size-tiered LSM-Tree, a tunable number $T$ of key duplicates are allowed to co-exist at a given level before their eventual merge into a larger level. Both compaction policies directly trade read cost against writes by altering the invariants of stored data.
Approximate member query (AMQ) filters, e.g. Bloom Filters (Bloom, 1970), allow memory to be spent to reduce the read cost of point queries by having a dedicated filter (or part of a shared filter (Conway et al., 2023)) devoted to each location where a key might reside. AMQ filters offer a tuning parameter $\alpha$, the false-positive rate, which trades size against accuracy. [^1] A filter must always answer accurately if a query key is not in its set, but it may give a false positive result at rate $\alpha$. The smaller $\alpha$ gets, the more bits per key are required to store the filter. To answer a point query for key $k$, the filters for all locations $S$ where $k$ might reside are consulted before performing any I/O. This rules out, with probability $\prod_{i\in S}{\left(1-\alpha_{i}\right)}$, locations where the query will fail, preventing excess reads.
If filters are used to maintain a target read cost in a size-tiered LSM-tree (with $T$ tiers per level), the total size of the filters must scale in proportion to $T$, since the total number of false-positives for a query is the sum of false-positives for all filters that pertain to a given key (Dayan et al., 2018). Therefore the $T$ parameter in a size-tiered LSM, together with $\alpha$, can be thought of as representing a trade-off between memory and write amplification. The problem with this mechanism is that the future read amplification one trades to achieve better write performance is reflected in the on-disk structure, and thus the memory cost for filters to compensate must be paid during the entire readable lifetime of write-optimized updates. Note that this is true even if the filter type (e.g. Quotient Maplets (Conway et al., 2023)) supports cache-friendly paging of partial filters into main memory, since the filter bit rate must nevertheless be increased to maintain a target query cost, regardless of how much of a filter is cached. In essence, write-optimization in the past generates a read/memory trade-off penalty in the future in such designs. We refer to this effect as the residual write-optimization penalty.
Hybrid-Optimized (Bϵ-trees). Bϵ-trees (Bender et al., 2015; Brodal and Fagerberg, 2003) are a tunable external memory dictionary, able to trade reads against writes. The basic B-tree structure is used as the foundation for a Bϵ-tree. The $\epsilon$ parameter varies the portion of a node page used for pivots, which optimize for reading, and update buffering, which optimizes for writing. At either extreme of their tuning range $\epsilon\in{0..1}$, Bϵ-trees simply degrade into the equivalent of a B-tree or LSM-tree. At $\epsilon=\frac{1}{2}$, the branching factor is $\sqrt{B}$, making tree height twice that of an equivalent B-tree. In this configuration, updates can be flushed down one level at a time from the root in batches of size $\Theta(\sqrt{B})$, dramatically lowering the worst-case write cost and providing a good balanced tuning point.
Size-Tiered Bϵ-trees, used by SplinterDB (Conway et al., 2020; Conway et al., 2023), are a recent re-imagining of the Bϵ-tree designed to close the ”bandwidth gap” between the capabilities of fast NVMe SSDs and traditional key/value storage engine designs. The main structural modification in STBϵ-trees is to add level of indirection between node pages and buffered data (Figure 2, 4 ). This change is analogous to the transformation between B-trees and B+-trees, and gives SplinterDB similar advantages in terms of being able to cache interior nodes for reading.
@@ 2.1. Current Approaches to Trade-off Tuning
SplinterDB adopts the strategy of selecting a single tuning point which attempts to offer excellent read and write efficiency by structuring per-node buffer data pages as a single-level, size-tiered LSM-tree-like structure, where each sorted run (or branch) is stored as a statically built B+-tree. Quotient Maplets, a variant of Quotient Filters, are used in this design to route point queries to the branches/tiers which may contain a given key.
While benchmark performance data (YCSB) shows this to be a good compromise for point-query workloads on modern hardware, it nonetheless represents a trade-off between false positives leading to increased read amplification on the one hand, and the memory cost of maintaining maplets with sufficient bits per key to compensate for the number of possible redundant copies of a key stored at multiple tiers. The downside of this approach is felt most acutely in short (¡100 key) range scans, which must merge/compact a potentially large number of tiers to determine sequential keys.
@@ 2.2. Towards Dynamic Tunability: Key Ideas
In this work, we propose a dynamic, unbiased approach that achieves desired read and write performance on demand while avoiding the costly re-tuning or data movement required in existing, statically-tuned systems.
Key Idea 0: Bϵ+-trees. A baseline requirement for dynamic optimization is an unbiased data structure with sufficient flexibility to achieve both high write and read performance. To achieve this property, we identify and build upon what we consider to be the core insight behind SplinterDB’s STBϵ-trees: moving buffered updates in a Bϵ-tree out of the node pages and into pages referenced from nodes (Figure 2, 3 $\Rightarrow$ 4 ). Following the observation that this transformation is quite similar to the difference between B-trees and B+-trees (Figure 2, 1 $\Rightarrow$ 2 ), we coin the term Bϵ+-tree to describe this new class of dictionaries.
Moving update data out of node pages improves flexibility by offering potential advantages for both read- and write-optimization in Bϵ+-trees. Read efficiency benefits from smaller nodes being more easily cached; likewise, queries benefit from reduced tree height as a result of a higher branching factor, since the portion of a node devoted to pivots can be larger without sacrificing buffer capacity. In terms of write efficiency, being able to buffer more data per node allows updates to be moved down the tree in larger batches, allowing more work to be achieved per write during flushes. The identification of Bϵ+-trees opens the door for research into a set of important design dimensions around how the data within a node buffer is structured. For example, instead of the STBϵ-tree’s flat structure of references to static B+-trees with size-tiered compaction policy, adopting a more greedy (level-tiered) compaction policy across levels within a node shifts away from a write-optimized focus towards a more unbiased one. A key contribution of this work is to begin exploring some of this space of possibilities to better understand their potential.
Key Idea 1: TurtleTrees. With Bϵ+-trees as a foundation, we introduce a novel data structure that leverages an unbiased, level-tiered node buffer structure and flush policy to enable dynamic read and write optimization. We name this new data structure the TurtleTree, shown in Figure 3. TurtleTrees are an instantiation of Bϵ+-trees in which buffer c -referenced data segments d are stored as flat leaf-like pages organized into a small, fixed capacity level-tiered LSM-tree per node a . Unlike size-tiered Bϵ+-trees, which are biased towards improving write performance at the cost of reads or log structured B-trees, which are biased towards improving reads over writes, the level-tiered TurtleTree sits at a middle point between these two poles. Like LSM-trees and other Bϵ-tree variants, TurtleTrees permit updates for a given key to be stored in multiple places (though keys must be unique within a given buffer level). By utilizing a level-tiered compaction rather than size-tiered, the number of potential locations is reduced, accelerating reads. TurtleTrees can also still use the various memory/read trade-off mechanisms discussed earlier, such as page caching and AMQ filtering, to further reduce read cost. In fact, we will see in this work that TurtleTrees can achieve read performance that is on par with or exceeds current state of the art read-optimized key/value stores.
@@
Setting aside reads, the question remains as to how to restore the write efficiency we lost by changing the intra-node compaction policy to level-tiered. We present a mechanism which achieves this in the next key idea (2). In preparation, we highlight a unique characteristic of TurtleTrees. By building a level-tiered structure within a Bϵ+-tree, we can bound the I/O cost of writing batches of modified data pages from the TurtleTree in-memory to disk, avoiding the write amplification penalty of traditional level-tiered LSM-trees. In fact, by setting the number of data pages d referenced by the buffer c roughly equal to the number of pivots b  [^2] The diagram in Figure 3 is conceptual; the exact ratios of pivots to buffer segments, and their relative sizes, are not accurate/to-scale., we guarantee that for each batch of pages inserted in the TurtleTree, we can flush an equal-sized batch to a lower level, bounding the growth of write amplification. This fine-grained control over how writes propagate down a TurtleTree allows us to choose exactly how many pages to flush and when to flush them.
Key Idea 2: Write Optimization via Checkpoint Distance. Given the unbiased TurtleTree structure, we can now introduce checkpoint distance, a tuning knob that improves write performance in exchange for memory without affecting future performance via residual optimization penalties. When receiving a batch of updates, TurtleKV applies the batch to a TurtleTree in-memory and can immediately write the modified pages to disk, creating a checkpoint. However, it is also possible to defer this write, buffering the changes from multiple batches within the TurtleTree in-memory before creating the corresponding checkpoint on disk. We refer to the number of batches $\chi$ buffered in-memory as the checkpoint distance. By modifying $\chi$ at runtime, it is possible to trade write performance for memory in TurtleKV; specifically, the number of writes required per key update is reduced by $O\left(\log{\chi}\right)$! This reduction is independent of the total size $N$ of stored data, making it a viable tuning knob regardless of the on-disk size or structure.
To understand why this is the case, consider the lifetime of one of the buffer data pages d shown in Figure 3. As new batches are applied to a node, the merging/compaction process causes data pages at higher levels to be frequently replaced with merged data and written to lower levels. This means that the closer a buffer segment’s level is to the root, the shorter its effective lifetime. By buffering batches in the TurtleTree, these short-lived data pages naturally expire and no longer need to be written to disk. Leveraging this property, we arrive at a key insight: by spending memory to defer writing these short-lived pages, we can avoid unnecessary writes and reduce write amplification without changing the stored structure of the eventual checkpoint.
Prior work such as SplinterDB (Conway et al., 2020; Conway et al., 2023) or TreeLine (Yu et al., 2022) permanently spend memory by storing additional filters or histograms to index into their data pages, incurring a residual optimization penalty each time they want to optimize performance. In contrast, checkpoint distance leverages the TurtleTree structure to improve write performance with no additional memory cost beyond the tree itself and its internal metadata. This means that checkpoint distance can be adjusted up or down to control write amplification without incurring any residual optimization penalty.
Key Idea 3: On-Demand Optimization. Instead of one-time, static optimization, TurtleKV leverages the TurtleTree structure to optimize both reads and writes on-demand, creating a general-purpose KV engine that negates the need for expensive data movement present in current Extract-Transform-Load (ETL) pipelines. TurtleTrees offer a tunable external memory dictionary which can navigate the RUM trilemma space using only read vs memory and write vs memory tuning mechanisms. We take advantage of this property by using checkpoint distance as a tuning knob to dynamically determine how much of the total available memory to spend on optimizing writes or reads. When the checkpoint distance is larger, more memory is spent buffering batches within TurtleTrees, optimizing writes, while when the checkpoint distance is smaller, more memory is dedicated in the page cache for data pages, optimizing reads. Since the tuning mechanism trades reads and writes against memory rather than reads vs writes directly, adjusting the memory allocation does not affect the underlying data structure. Thus, TurtleKV can dynamically adjust the memory allocation without permanently affecting either read or write performance, allowing for on-demand optimization of reads or writes depending on application requirements.
In systems whose tuning mechanisms trade reads against writes directly, such as compaction policy in LSM-tree storage engines, the cost of re-tuning is high because it involves changing the stored structure of data to target a new trade-off point. Historically, applications have implemented this kind of ”tuning by moving” strategy through the use of multiple data stores and Extract-Transform-Load (ETL) pipelines, which transfer data between storage systems to optimize it for various application requirements. The fact that ETL and multi-stores have been quite popular despite high cost and increased system complexity speaks to the need for data management systems which can re-tune on demand to support changing workloads. In contrast to expensive data movement, TurtleKV can simply spend memory whenever it is advantageous to do so to accelerate read- or write-heavy workloads: pay now, not later.
3. TurtleTrees & Checkpoint Distance
[…]
Compatibility with other mechanisms. We focus on $\chi$ as the main parameter to improve TurtleTrees’ performance. TurtleTrees are also compatible with other mechanisms to increase read performance: increasing page-cache size and read-filter bitrate. In this work, we assume static sizes for these additional parameters to isolate the effect of TurtleTrees and $\chi$.
@@ 3.1. TurtleTree Structure
TurtleTrees are Bϵ+-trees ($\S$2.2) whose per-node update buffer is organized as a fixed-capacity, level-tiered LSM-tree with fanout F=2F=2.333That is, level to level growth factor; not to be confused with the branching factor of the node, or number of pivots, which we denote as $\rho$. The structure of TurtleTree node and leaf pages is shown in Figure 4.
Leaf pages contain a sorted list of key/value pairs. As is typical of B+-trees, leaves are sized several orders of magnitude larger than interior nodes. [^4] A typical configuration might be 4KB node pages and 16MB leaf pages. To ensure fast in-memory queries (both point and range/scan), TurtleKV leaf pages contain a cache-oblivious, byte-addressed binary trie index. This index contains every 16th or 32th key from the leaf, depending on available space. The trie is made cache-oblivious by laying out nodes in Van Emde Boas order (Bender et al., 2007) to maximize locality of reference along any given query path.
Each (interior) node is split into roughly equal halves, with one half devoted to pivots (dark gray) and the other to the update buffer (yellow). Within a node page, the update buffer region contains pointers to segments (the blue boxes with rounded sides), which are stored as leaf pages with the addition of tombstone records for deleted keys. The segments in a buffer are stratified into levels of exponentially increasing size. Keys are required to be unique and sorted within a level. The first level contains at most a single segment, the second level at most two, the third level four, then eight, etc. All levels start out vacant and can be in either a vacant or occupied state. For a given number of pivots $\rho$ in a node, the number of levels is set to $\lceil\log_{2}{\rho}\rceil$, and the total number of segments in all levels of the buffer bounded to at most $\rho-1$.
Figure 4. Details of the internal structure of TurtleKV’s TurtleTree implementation.
@@
3.2. TurtleTree Batch Updates
@@
Figure 5. Node-Buffer Insertion of Batch Updates in TurtleTrees.
@@
3.3. Checkpoint Distance

### v2

@@ 2.2.2. RocksDB
(a) SplinterDB
(b) TurtleKV
[…]
Figure 3b shows the result of repeatedly running a uniform random insertion workload against an empty RocksDB database, while increasing $\overline{M_{w}}$ from 16mb up to 4gb. When 16mb $\leq\overline{M_{w}}\leq$ 1gb, we see the predicted logarithmic reduction in write amplification (x-axis is log scale). However, for MemTable/SSTable sizes over 1gb, the write optimization effect flattens out, since the inserted/flushed data never escapes $L_{0}$ before the workload completes [^8] Key ranges of SSTable files in $L_{0}$ are allowed to overlap, so as not to stall MemTable flushes; at least 4 files must be present in $L_{0}$ to trigger a compaction to $L_{1}$.. Unfortunately this reduction in write volume does not translate into lower average insert latencies beyond RocksDB’s default value of $\overline{M_{w}}$ =64mb. This underscores the necessity of optimizing main-memory costs (CPU, memory bandwidth, and internal synchronization bottlenecks) if one targets I/O reduction as a tuning mechanism.
2.2.3. SplinterDB
@@
SplinterDB largely adopts the strategy of selecting a single tuning point which attempts to offer excellent read and write efficiency by structuring per-node buffer data pages as a single-level, size-tiered LSM-tree-like structure, where each sorted run (or branch) is stored as a statically built B+-tree. Quotient Maplets, a variant of Quotient Filters (pandeyVectorQuotientFilters2021a), are used in this design to route point queries to the branches/tiers which may contain a given key. SplinterDB’s STBϵ-trees have configurable branching factor, which is tied to the number of tiers per level $T$. Quotient filter bit rate may also be tuned, and thus STBϵ-trees theoretically support a degree of write-memory tuning via filter arbitrage.
[…]
2.2.4. TurtleKV
For comparison, we show TurtleKV’s results for both N=100M write-buffer size scaling and N=500M cache size-scaling workloads in Figures 3c and 4b respectfully. These results show that TurtleKV’s primary data structure, the TurtleTree, efficiently reduces write cost at a rate of $O(\log(M))$, where $M$ is the memory spent on write-optimization, like level-tiered LSM-tree-based RocksDB. However, because of the system design ideas elaborated in the following sections, TurtleKV is able to turn WAF reduction into higher performance over a very wide tuning parameter range, similar to WiredTiger. When configured for write optimization, TurtleKV keeps pace with SOTA engine SplinterDB, but because TurtleKV’s data-parallel strategy for in-memory compactions is able to achieve even load balancing across processors, its scaling curve is much smoother than SplinterDB’s (Figure 4).
@@ 2.3. Towards Dynamic Tunability: Key Ideas
Table 1 expresses the three high-level goals which drive the design of TurtleKV. In this section we provide a brief sketch of the key ideas which address each goal. These ideas will be elaborated in the following sections.
(a) Logical Structure
(b) Page Layout
@@
2.3.1. Data Structure With Efficient WM Trade-off
We start with the STBϵ-tree, observing that its desirable properties derive from the trunk/branch separation, whereas the properties we would like to change derive from their compaction and flush policies. We coin the term Bϵ+-tree to describe the class of dictionaries which retain the trunk/branch separation, but which leave open the questions of flush/compaction policy and how (if at all) branches are structured within their referent trunk node(s) (fig. 2). Our first key idea is to instantiate the Bϵ+-tree with a balanced, level-tiered compaction policy which organizes branches into sub-levels within a trunk node, as shown in Figure 5a. This data structure we name the TurtleTree.
2.3.2. System Design With No In-Memory Bottlenecks
TurtleKV adopts a design similar to LSM-tree-based systems like RocksDB, with a write-ahead log (WAL), in-memory component (MemTable), and log-structured on-disk component (Checkpoint, a TurtleTree). The primary bottleneck in this design comes from the $\log(N)$ amplification when merging MemTables into the current Checkpoint. We observe that the most CPU-intensive part of this process is the key comparisons for level-to-level merging, and the most memory-bandwidth-intensive part is the copying of key-value data to serialize compacted data at its new location. Our second key idea is to use data-parallel algorithms for level merging, and fork/join style multi-threaded leaf page serialization for Checkpoint writes.
2.3.3. First-Class Write-Memory (WM) Tuning Knob
Inspired by our RocksDB case study, we identify MemTable size $\overline{M_{w}}$ as our WM tuning knob. This defines the distance, in total bytes of updates, between Checkpoints. To ensure that increasing checkpoint distance does not degrade in-memory query performance, we build a single large MemTable between checkpoint updates. To keep tail latencies under control, and to avoid affecting the on-disk structure of checkpoints, we break finalized MemTables into leaf-page sized batches by doing a key-order scan, applying the batches to the Checkpoint TurtleTree one at a time in-cache until the MemTable is drained. This is our third key idea, which we call checkpoint distance tuning with Big MemTables.
3. TurtleTrees
In this section we describe the TurtleTree data structure, its core algorithms (batch update, get, and scan), and how it leverages checkpoint distance as a WM trade-off tuning mechanism.
@@ 3.1. TurtleTree Structure
TurtleTrees are Bϵ+-trees ($\S$2.3) whose per-node update buffer is organized into a fixed number of levels as shown in Figure. 5a. The structure of TurtleTree node and leaf pages is shown in Figure. 5b.
3.1.1. Leaf Pages
Leaf pages contain a sorted list of key/value pairs. To achieve high I/O utilization during updates and scans, leaf pages are sized several orders of magnitude larger than interior nodes. [^9] The configuration we use is 4KB node pages and 32MB leaf pages. In order to guarantee low latencies during point queries, leaf pages are internally structured to allow small slices, or sharded views, to be cached and re-used to find queried keys. Leaves also include a cache-oblivious, byte-addressed binary trie index at the front of the page. This index contains every $k^{th}$ key from the leaf, $k\in\{1..32\}$. The trie is made cache-oblivious by laying out nodes in Van Emde Boas order (benderCacheobliviousStreamingBtrees2007a) to maximize locality of reference along any given query path, minimizing CPU cache misses for in-memory queries.
3.1.2. Node Pages
Each (interior) node is split into roughly equal halves, [^10] The intra-page regions in Figure 5b are not shown to scale. with one half devoted to pivots (dark gray) and the other to the update buffer (yellow). Within a node page, the update buffer region contains pointers to segments, which are stored as leaf pages with the addition of tombstone records for deleted keys. The segments in a buffer are stratified into levels of exponentially increasing size. Keys are required to be in-order and unique within a level. The first level contains at most a single segment, the second level at most two, the third level four, then eight, etc. All levels start out vacant and can be in either a vacant or occupied state. For a given number of pivots $\rho$ in a node, the number of levels is limited to $\lceil\log_{2}{\rho}\rceil$, and the total number of segments in all levels of the buffer limited to at most $\rho-1$.
@@ 3.1.2. Node Pages
3.2. TurtleTree Operations
3.2.1. Batch Updates
@@
3.2.2. Queries
Point and range queries are the same for TurtleTrees as for Bϵ-trees, with the addition of chasing pointers to segment page(s) in the buffer (instead of just reading its contents). We use the active pivots bit sets stored for each segment to narrow the search to only those which might contain the key. Further pruning of the set of candidate pages to search is done by querying the quotient filter associated with a given segment or leaf.
After finding a key in a segment, we compare its position in the leaf to the flushed upper bound for its corresponding pivot; if this position is less than the upper bound (i.e., it has been marked as flushed), we act as if the key is not present and continue the search.
Deletes are performed alongside Put operations as part of a batch update, by setting a key’s value to the special DELETED tombstone value. This value is preserved in leaves until the bottom of the tree is reached; at this point, the key is removed from the merged output.
3.3. TurtleTree Cost Analysis
3.3.1. Update I/O cost
Individual key updates travel from root to leaf of the tree as batch updates and flushes take place. Since there are $O(\log_{2}{\rho})$ levels within the buffer of each node, and the tree height is $O(\log_{\rho}{(N/L)})$, each key update is written $O(\log_{2}{(N/L)})$ $O(\log_{2}{\rho}\cdot\log_{\rho}{(N/L)})$ times, regardless of page size or branching factor $\rho$.  [^13] Note that in order to correctly analyze the cost of a key update, we need only consider its initial journey from root-to-leaf, since each time a key’s page is re-written after it reaches the bottom of the tree, we can amortize that cost against the incoming batch, which represents an equal number/size of keys ending their root-to-leaf journey. If we define the transfer size $B$ in units equal to a single key update, the I/O cost of a key update is $O(B^{-1}\cdot\log_{2}{(N/L)})$.
3.3.2. Query cost
Point queries (Get) differ in their I/O cost depending on whether or not the key is found. Here we give a worst-case bound for the general case; in practice this can be dramatically improved through the addition of per-leaf filters, as described in $\S$4.1.2. The cost analysis here also depends on the specific I/O model used; if we are counting total page loads, as in the Disk Access Model (DAM), then a single-key (point) query must load all node pages and buffer segments along a single root-to-leaf path: $O(\log_{\rho}{(N/L)}+\frac{L}{B}\cdot\log_{2}{(N/L)})$ or just $O(\frac{L}{B}\cdot\log_{2}{\frac{N}{L}})$.
Like other B-tree variants, range queries of length $k$ on Turtle Trees cost a point query (to find the starting key) plus $O(k/B)$ IOPs. There is also an in-memory, per-key cost of $\log_{2}{(\log_{2}{(N/L)})}$ to merge sorted runs, which is the same as a level-tiered LSM-tree. This per-key cost expression is derived via the use of a binary min-heap to select which of the $\log_{2}{(N/L)}$ levels of the tree contains the least-ordered key.
3.3.3. Write-Memory Trade-off Efficiency
The WM tuning knob for TurtleTrees is a parameter chi ($\chi$), for checkpoint distance, which determines the number of batches applied to a TurtleTree in-memory before flushing changed pages to storage as a new checkpoint. At $\chi=1$, every time we perform a batch update on the tree, all modified pages must be written back out. At $\chi=2$, we defer the writing of new/modified pages until two batches have been applied, etc. It may not be immediately obvious to the reader that $\chi$ should have any effect on the write cost of key updates. To see that it does, let us revisit the batch insertion example illustrated by Figure 6, considering the span of time (as measured by batch number) that each page is live, shown in Figure 7. The pages which appear in durable storage at a given time step are those whose lifespan overlaps a given column in the middle section (Lifetime) of the diagram.